5.30. UITULU eur 5.39. Evaluate z dz when : >0 and C is the circle Izl = 3. 2 Ti I (z2 + 1)
Answer #1

het het I= Il ezt az ? 201. (z4412 C: 121=3 (22+1,8 =0 2+1=0 z=ti atz = ti is of poke of order 2. Now residue at 2= 2 = (a $(Therefore I T I = ani - - I Perdz ni Jo (22 +12 sum of cum an are residue at z=i&i. at 7=;8 - 1 c (2 I= – tcost sint - t Cost